OpenStudy (anonymous):

log[2](b)8=v and log[2](b)5=h log[2](b)√(8/125)

OpenStudy (anonymous):

how to write the expression in terms of v and h

OpenStudy (anonymous):

you need to use the calculator

OpenStudy (anonymous):

i can't i'm trying to express it in terms of v and h

OpenStudy (anonymous):

im not sure then sorry

OpenStudy (anonymous):

ok thanks

zepdrix (zepdrix):

I don't understand your logs the way you wrote them. Is 2 the base? Or is the base 2b? What is that 8 doing there? Is it a base of 2, with b to the 8th power inside the log?

OpenStudy (anonymous):

no b is the base sorry...another website added the 2 it is logb8=v and logb5=h and im trying to find logbsqrt(8/125)

OpenStudy (anonymous):

anytime it says(2) it means the next number is a base

zepdrix (zepdrix):

oh i see :)

OpenStudy (jdoe0001):

\(\bf log_2b^8=v\\ log_25^b=h\\ log_2\left(b\sqrt{\cfrac{8}{25}}\right) \ \ \ ?\)

OpenStudy (jdoe0001):

shoot I even mixed up numbers, one sec

OpenStudy (jdoe0001):

\(\bf log_2b^8=v\\ log_2b^5=h\\ log_2\left(b\sqrt{\cfrac{8}{25}}\right) \ \ \ ?\)

OpenStudy (jdoe0001):

is that what you meant?

OpenStudy (anonymous):

no, the base is a b...forget the 2

OpenStudy (jdoe0001):

hmm, how about you use the whiteboard? :)

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

i dont even know what that it

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

or post a screenshot

zepdrix (zepdrix):

\[\Large \log_b8=v \qquad \log_b5=h\]And we need to write this in terms of v and h?\[\Large \log_b \sqrt{\frac{8}{125}}\]

OpenStudy (anonymous):

yes zepdrix

zepdrix (zepdrix):

We'll use properties of exponents and logs to mess with this thing, and get it looking the way we want.\[\Large \log_b \sqrt{\frac{8}{125}} \qquad=\qquad \log_b\frac{\sqrt 8}{\sqrt {125}}\qquad=\qquad \log_b\frac{8^{1/2}}{5^{3/2}}\]Confused by any of that so far?

OpenStudy (anonymous):

nope that seems good so far

OpenStudy (jdoe0001):

keep in mind that \(\bf \Large {\log_b8=v \implies b^v = 8\\ \log_b5=h \implies b^h = 5}\)

zepdrix (zepdrix):

A property of logs tells us,\[\large \log(\frac{a}{b})=\log(a)-\log(b)\] So let's apply this to our problem.\[\Large \log_b\frac{8^{1/2}}{5^{3/2}} \qquad=\qquad \log_b8^{1/2}-\log_b5^{3/2}\]

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

Another important property of logs,\[\Large \log(a^c)=c\cdot \log(a)\] Let's use this rule to move our fractions in front as coefficients.

zepdrix (zepdrix):

\[\Large \frac{1}{2}\log_b8-\frac{3}{2}\log_b5\]Hmm we can probably plug in our v and h from here :)

zepdrix (zepdrix):

\[\Large \frac{1}{2}\color{#FF11AA}{\log_b8}-\frac{3}{2}\color{#3366CF}{\log_b5}\] And let's remember our v and h equations they gave us,\[\Large \color{#FF11AA}{\log_b8=v} \qquad \color{#3366CF}{\log_b5=h}\]

OpenStudy (anonymous):

1/2v-3/2h?

zepdrix (zepdrix):

Yah that sounds right! :) Did you get confused by what I did with the 125, turning it into a 5? That's a bit of a tricky step.

OpenStudy (anonymous):

no you didnt thanks so much

OpenStudy (anonymous):

the statue of liberty is approximately 305 ft tall. if the angle of elevation of a ship is 28.6 degrees, how far, to the nearest foot is the ship from the statues base any clue with this one

zepdrix (zepdrix):

Mmmm did I draw that angle in the correct location? Sometimes I mess that up. Angle of Elevation.. hmm